# DP Technology *

### v1

@@ 1 Introduction
Current approaches, both pipeline-based [livathinos2025docling, cui2025paddleocr, wang2024mineru, mathpix] and VLM-based [blecher2023nougat, wei2024got, nassar2025smoldocling, li2025monkeyocr, ocrflux, rednote_hilab_dots_ocr_2025, mistral_ocr], face three major challenges. First, they are computationally inefficient, making it prohibitively costly to parse tens of millions of documents at scale. Second, their performance on non-textual modalities such as formulas, tables, charts, and chemical structures remains limited, with low accuracy and poor robustness. Third, the complex and heterogeneous layouts of scientific and patent documents often lead to unreliable segmentation and structural analysis, further degrading downstream usability. Pipeline methods offer higher throughput but poor generalization, while VLM methods generalize better but suffer from low efficiency and limited extensibility. Collectively, these limitations hinder the construction of large-scale, high-quality knowledge bases required for both academic research and industrial innovation.
@@ 2.1 Overall Framework
Group-based Layout Detection: The layout detection model locates each semantic block in the PDF pages and identifies its semantic category. Unlike conventional approaches [cui2025paddleocr, wang2024mineru], our group-based layout detection model recognizes naturally paired semantic components—such as image–caption, table–title, and molecule–identifier—preserving visual–semantic associations that are critical for structured information extraction and accurate reading-order construction. Finally, different semantic blocks are rendered at dynamic resolutions and forwarded to designated models for image-to-text processing. We employ a greedy batch-stacking strategy for rendering page images, performing layout recognition, and dispatching semantic groups to downstream microservices. As soon as a batch accumulates in the queue, it is immediately processed by subsequent modules, allowing the latency of this stage to be almost entirely masked by other processing steps.
@@ 2.2 Group-based Layout Analysis
Layout detection is a foundational prerequisite for all subsequent document analysis tasks and represents a pivotal component within document parsing systems. Its efficacy ultimately establishes the performance ceiling for the entire system. Acknowledging the profound diversity and complexity of scientific document layouts, along with the critical importance of grouping relationships among heterogeneous semantic elements, we introduce a layout understanding paradigm that diverges significantly from conventional approaches [cui2025paddleocr, zhao2024doclayout, pfitzmann2022doclaynet, livathinos2025docling]: a group-based tree-structured layout representation.
@@
Image3 Figure / Multi-modal Text Image Caption
@@
Considering the wide variety of authoring and rendering styles across layouts and modalities, we employ a large-scale dataset of PDF pages for training. We built an in-house layout detection dataset containing 500k pages. Of these, 220k pages are carefully human annotated with group-based layout labels from a diverse corpus. The database primarily consists of scientific journal data and patent data from various patent offices. It also includes preprints, books, and other types of data across different fields, spanning a total of 85 languages. For more details, refer to Section 4. The remaining pages are synthetic data used for pretraining. Our experiments show that when high-quality, large-scale real data is available, pretraining with existing synthetic datasets, such as DocSynth300K [zhao2024doclayout], can be counterproductive. This is due to their low-fidelity rendering, stylistic deviations from authentic documents, and limited diversity imposed by manually defined generation rules, which fail to capture the complexity of real-world layouts. To address this, we generate synthetic samples from real layouts using controlled modifications—including element merging, spatial perturbations, and semantic content substitutions—allowing the model to effectively leverage synthetic data for pretraining.
To meet real-time processing requirements, our selection of a backbone architecture for layout detection prioritizes high efficiency. A survey of contemporary object detection methods reveals that models such as RT-DETRv2 [rtdetr], YOLOv12 [tian2025yolov12], and D-FINE [peng2024dfine] can exhibit training instability in our layout detection task. Meanwhile, the extensive modifications made to YOLOv8 [jocher2023ultralytics] in DocLayout-YOLO [zhao2024doclayout] yield only marginal performance improvements in our setting. Consequently, after balancing performance with inference speed, we select a modified DETR-based architecture as the backbone for our group-based layout detection model.
The choice of input resolution presents another critical trade-off between processing speed and detection accuracy. While prior research [steiner2024paligemma] suggests that a resolution of $896\times 896$ suffices for most documents, our empirical evaluations demonstrate its insufficiency for achieving consistently high performance. Therefore, to accommodate diverse document aspect ratios and strike a superior balance between speed and accuracy, we adopt an input resolution of $1024\times 768$.
@@ 2.3 OCR
Uni-Parser operates in two complementary OCR modes: extraction and recognition. When textual content can be directly extracted from the PDF in the first stage, the system bypasses OCR and utilizes the extracted text directly. Otherwise, or when explicitly requested by the user, Uni-Parser performs recognition using OCR model, such as the PP-OCRv5 [cui2025paddleocr] series. The default OCR backbone supports a broad character set encompassing Simplified and Traditional Chinese, English, Japanese, and a wide range of special symbols. For other languages, the system automatically invokes the corresponding language-specific PP-OCRv5 model. For more complex cases—such as multilingual mixtures or the presence of numerous out-of-vocabulary OCR symbols—we additionally support switching to PaddleOCR-VL [cui2025paddleocrvlboostingmultilingualdocument] in high-accuracy mode, trading speed for improved robustness in challenging scenarios. In practice, additional preprocessing is applied in certain scenarios. For example, when text is embedded within tables, Uni-Parser employs a text orientation detector to ensure accurate recognition.
@@ 2.4 Table Structure Recognition
We adopt an efficient table structure recognition model, SLANet [li2022ppslanet], trained on a corpus of one million tables that integrates a cleaned version of PubTabNet [zhong2019image], SynthTabNet [nassar2022tableformer], and our in-house synthetic dataset of line-based tables enriched with multimodal content. The synthetic dataset covers both bordered and borderless formats and incorporates diverse elements, including molecule structures, charts, images, formulas, and text. This comprehensive training strategy allows SLANet to achieve strong generalization and superior performance on complex real-world table layouts.
@@ 2.5 Formula Recognition
The system integrates a mathematical formula recognition module fine-tuned from PP-Formula [cui2025paddleocr]. The module converts mathematical expressions and chemical equations into LaTeX sequences in an end-to-end manner. It handles both standalone formulas and inline formulas or equations within paragraph text. For standalone formulas, the layout stage groups the expression with its reference number to form a structured representation.
@@ 2.6 Chemical Expression Recognition
Chemical structures are a fundamental modality in scientific literature and patents, particularly in chemistry, pharmaceuticals, biology, and materials. Uni-Parser adopts the end-to-end OCSR architecture MolParser [fang2024molparser] for molecular recognition, which translates molecular images into an Extended SMILES (E-SMILES) representation. We further introduce MolParser 1.5, which extends MolParser with an expanded pretraining corpus. In addition to the synthetic MolParser-7M dataset, we add a real-world dataset of 10 million images pseudo-labeled via cross-validation using multiple MolParser models and a fine-tuned MolScribe [qian2023molscribe]. With a larger proportion of in-the-wild chiral molecules and Markush structures, MolParser 1.5 yields a more balanced and comprehensive pretraining dataset, and after fine-tuning on the MolParser-SFT dataset, it produces a more robust end-to-end OCSR model.
@@ 2.7 Chart and Scientific Figure Understanding
For chart understanding, we fine-tuned a Qwen-2.5-VL-3B [bai2025qwen2] model on a dataset consisting of 500k generated samples, 300k open-source samples, and 170k real-world charts. This training significantly enhances the model’s ability to convert charts into tables, with a particular focus on multi-subplot charts, which frequently appear in scientific literature.
To support accurate scientific figure captioning, we curated a large-scale, high-quality dataset of 4 million samples. The core of this dataset consists of 3 million scientific images meticulously collected from high-impact publications. To enhance caption quality, we leveraged multiple multi-modal large language models (MLLMs), including but not limited to GPT-5 [openai2025gpt5] and Gemini 2.5 [deepmind2025gemini25pro], to rewrite the original captions by integrating visual content with raw captions and relevant contextual information from the associated papers. The dataset is further augmented with additional sources, including ChemPile [mirza2025chempile], molecular descriptions from PubChem, Electron Microscopy image captions [wang2025uniem3muniversalelectronmicrograph], and other experimental characterization datasets. Leveraging this comprehensive dataset, we fine-tuned a Qwen-2.5-VL-3B model, which named as SciParser, enabling it to generate precise, context-aware captions for a diverse range of scientific figures.
@@ 3.2 Deployment Scaling
The Uni-Parser system has been scaled to a cluster of 240 NVIDIA L40 (48GB) GPUs, with each GPU allocated 22 CPU cores and 90GB of memory. Using this configuration, the system parses more than 16 million documents in less than 6 days (in Uni-Parser fast mode). These results demonstrate that the Uni-Parser infrastructure is highly efficient, economically scalable, and robustly stable under large-scale processing workloads.
@@ 4.1 Uni-Parser Data Engine
Step 2: Active Learning on Real Documents. After the model is equipped with basic capabilities using synthetic data, we leverage a large corpus of real-world documents to iteratively enhance its performance through active learning. Our in-house training datasets are curated from:
  • 170 million pages of scientific literature spanning multiple disciplines;
  • 140 million pages of patent documents from patent offices worldwide;
  • 20 million pages of books, reports, and other documents sourced from Fine-PDFs [kydlicek2025finepdfs].
An active learning data flywheel, similar to the one used in MolParser [fang2024molparser], is employed to select the most informative samples for annotation. This is combined with a curriculum learning strategy that progressively trains the models on increasingly complex examples.
@@ 5 Performance
5.1 Construction of Uni-Parser Benchmark
Uni-Parser benchmark dataset targets layout detection and semantic contents recognition in scientific documents, comprising 150 PDFs spanning international patents and research articles. The selection emphasizes diversity in structure, subject, domain, and language. The distribution of this benchmark is shown in Table  2.
@@ 5.2 Chemical Structure Parsing
We evaluate Uni-Parser on both the Uni-Parser Benchmark and the recent third-party BioVista benchmark [yan2025biominer]. As shown in Table 3(a), our group-based layout detection model, Uni-Parser-LD, outperforms our previous state-of-the-art molecular detection model, MolDet [fang2024molparser], while also establishing correspondences between molecules and their identifiers—a capability absent in MolDet. Table 3(b) shows that MolParser 1.5 consistently surpasses its predecessor across all benchmark subsets, delivering notable improvements in both accuracy and inference speed over mainstream open-source OCSR methods. Moreover, our approach remains highly competitive against costly commercial models [MolVision_2025, xiong2023alphaextractor]. Representative case studies are provided in Appendix B.1.
@@
MolDet-Doc-L[fang2024molparser] 0.983 0.919
MolDet-General-L[fang2024molparser] 0.974 0.815
@@
MolDet-Doc-L[fang2024molparser] 0.961 0.871
MolDet-General-L[fang2024molparser] 0.945 0.815
BioMiner[yan2025biominer] 0.929 -
MolMiner[molminer] 0.899 -
@@
MolParser 1.0[fang2024molparser] 0.953 0.676 0.664 0.800
MolScribe[qian2023molscribe] 0.617 0.274 0.168 0.417
@@
MolParser 1.0[fang2024molparser] 0.669 0.352 0.733 0.703
MolMiner[molminer] 0.774 0.497 0.185 0.507
MolScribe[qian2023molscribe] 0.703 0.481 0.156 0.455
MolNexTR[chen2024molnextr] 0.695 0.419 0.045 0.401
DECIMER[rajan2023decimer] 0.545 0.326 0.000 0.298
@@
Table 4: Comparison of PDF parsing frameworks with OCSR support on a tiny test set, provided purely as a qualitative analysis rather than a quantitative evaluation.
@@
MathPix [mathpix] 2025-08-27 100% 75.9% 59.6% - 66.1 s
MinerU.Chem [MinerU.Chem_2025] 2025-09-17 66.7% 63.1% 22.7% - $\sim$7 min

### v2

@@ 1 Introduction
Current approaches, both pipeline-based [18, 7, 43, 20] and VLM-based [2, 45, 25, 15, 26, 36, 23], face three major challenges. First, they are computationally inefficient, making it prohibitively costly to parse tens of millions of documents at scale. Second, their performance on non-textual modalities such as formulas, tables, charts, and chemical structures remains limited, with low accuracy and poor robustness. Third, the complex and heterogeneous layouts of scientific and patent documents often lead to unreliable segmentation and structural analysis, further degrading downstream usability. Pipeline methods offer higher throughput but poor generalization, while VLM methods generalize better but suffer from low efficiency and limited extensibility. Collectively, these limitations hinder the construction of large-scale, high-quality knowledge bases required for both academic research and industrial innovation.
@@ 2.1 Overall Framework
Group-based Layout Detection: The layout detection model locates each semantic block in the PDF pages and identifies its semantic category. Unlike conventional approaches [7, 43], our group-based layout detection model recognizes naturally paired semantic components—such as image–caption, table–title, and molecule–identifier—preserving visual–semantic associations that are critical for structured information extraction and accurate reading-order construction. Finally, different semantic blocks are rendered at dynamic resolutions and forwarded to designated models for image-to-text processing. We employ a greedy batch-stacking strategy for rendering page images, performing layout recognition, and dispatching semantic groups to downstream microservices. As soon as a batch accumulates in the queue, it is immediately processed by subsequent modules, allowing the latency of this stage to be almost entirely masked by other processing steps.
@@ 2.2 Group-based Layout Analysis
Layout detection is a foundational prerequisite for all subsequent document analysis tasks and represents a pivotal component within document parsing systems. Its efficacy ultimately establishes the performance ceiling for the entire system. Acknowledging the profound diversity and complexity of scientific document layouts, along with the critical importance of grouping relationships among heterogeneous semantic elements, we introduce a layout understanding paradigm that diverges significantly from conventional approaches [7, 54, 32, 18]: a group-based tree-structured layout representation.
@@
Image3 Figure / Multi-modal Text -
@@
Considering the wide variety of authoring and rendering styles across layouts and modalities, we employ a large-scale dataset of PDF pages for training. We built an in-house layout detection dataset containing 500k pages. Of these, 220k pages are carefully human annotated with group-based layout labels from a diverse corpus. The database primarily consists of scientific journal data and patent data from various patent offices. It also includes preprints, books, and other types of data across different fields, spanning a total of 85 languages. For more details, refer to Section 4. The remaining pages are synthetic data used for pretraining. Our experiments show that when high-quality, large-scale real data is available, pretraining with existing synthetic datasets, such as DocSynth300K [54], can be counterproductive. This is due to their low-fidelity rendering, stylistic deviations from authentic documents, and limited diversity imposed by manually defined generation rules, which fail to capture the complexity of real-world layouts. To address this, we generate synthetic samples from real layouts using controlled modifications—including element merging, spatial perturbations, and semantic content substitutions—allowing the model to effectively leverage synthetic data for pretraining.
To meet real-time processing requirements, our selection of a backbone architecture for layout detection prioritizes high efficiency. A survey of contemporary object detection methods reveals that models such as RT-DETRv2 [53], YOLOv12 [41], and D-FINE [31] can exhibit training instability in our layout detection task. Meanwhile, the extensive modifications made to YOLOv8 [11] in DocLayout-YOLO [54] yield only marginal performance improvements in our setting. Consequently, after balancing performance with inference speed, we select a modified DETR-based architecture as the backbone for our group-based layout detection model.
The choice of input resolution presents another critical trade-off between processing speed and detection accuracy. While prior research [38] suggests that a resolution of $896\times 896$ suffices for most documents, our empirical evaluations demonstrate its insufficiency for achieving consistently high performance. Therefore, to accommodate diverse document aspect ratios and strike a superior balance between speed and accuracy, we adopt an input resolution of $1024\times 768$.
@@ 2.3 OCR
Uni-Parser operates in two complementary OCR modes: extraction and recognition. When textual content can be directly extracted from the PDF in the first stage, the system bypasses OCR and utilizes the extracted text directly. Otherwise, or when explicitly requested by the user, Uni-Parser performs recognition using OCR model, such as the PP-OCRv5 [7] series. The default OCR backbone supports a broad character set encompassing Simplified and Traditional Chinese, English, Japanese, and a wide range of special symbols. For other languages, the system automatically invokes the corresponding language-specific PP-OCRv5 model. For more complex cases—such as multilingual mixtures or documents containing numerous out-of-vocabulary OCR symbols—we additionally support switching to PaddleOCR-VL [6] in a high-quality mode, trading inference speed for improved robustness in challenging scenarios, with support for more than 109 languages (ISO 639). In practice, additional preprocessing is applied in certain scenarios. For example, when text is embedded within tables, Uni-Parser employs a text orientation detector to ensure accurate recognition.
@@ 2.4 Table Structure Recognition
We adopt an efficient table structure recognition model, SLANet [13], trained on a corpus of one million tables that integrates a cleaned version of PubTabNet [55], SynthTabNet [24], and our in-house synthetic dataset of line-based tables enriched with multimodal content. The synthetic dataset covers both bordered and borderless formats and incorporates diverse elements, including molecule structures, charts, images, formulas, and text. This comprehensive training strategy allows SLANet to achieve strong generalization and superior performance on complex real-world table layouts.
@@ 2.5 Formula Recognition
The system integrates a mathematical formula recognition module fine-tuned from PP-Formula [7]. The module converts mathematical expressions and chemical equations into LaTeX sequences in an end-to-end manner. It handles both standalone formulas and inline formulas or equations within paragraph text. For standalone formulas, the layout stage groups the expression with its reference number to form a structured representation.
@@ 2.6 Chemical Expression Recognition
Chemical structures are a fundamental modality in scientific literature and patents, particularly in chemistry, pharmaceuticals, biology, and materials. Uni-Parser adopts the end-to-end OCSR architecture MolParser [9] for molecular recognition, which translates molecular images into an Extended SMILES (E-SMILES) representation. We further introduce MolParser 1.5, which extends MolParser with an expanded pretraining corpus. In addition to the synthetic MolParser-7M dataset, we add a real-world dataset of 10 million images pseudo-labeled via cross-validation using multiple MolParser models and a fine-tuned MolScribe [34]. With a larger proportion of in-the-wild chiral molecules and Markush structures, MolParser 1.5 yields a more balanced and comprehensive pretraining dataset, and after fine-tuning on the MolParser-SFT dataset, it produces a more robust end-to-end OCSR model.
@@ 2.7 Chart and Scientific Figure Understanding
For chart understanding, we fine-tuned a Qwen-2.5-VL-3B [1] model on a dataset consisting of 500k generated samples, 300k open-source samples, and 170k real-world charts. This training significantly enhances the model’s ability to convert charts into tables, with a particular focus on multi-subplot charts, which frequently appear in scientific literature.
To support accurate scientific figure captioning, we curated a large-scale, high-quality dataset of 4 million samples. The core of this dataset consists of 3 million scientific images meticulously collected from high-impact publications. To enhance caption quality, we leveraged multiple multi-modal large language models (MLLMs), including but not limited to GPT-5 [28] and Gemini 2.5 [8], to rewrite the original captions by integrating visual content with raw captions and relevant contextual information from the associated papers. The dataset is further augmented with additional sources, including ChemPile [22], molecular descriptions from PubChem, Electron Microscopy image captions [47], and other experimental characterization datasets. Leveraging this comprehensive dataset, we fine-tuned a Qwen-2.5-VL-3B model, which named as SciParser, enabling it to generate precise, context-aware captions for a diverse range of scientific figures.
@@ 3.2 Deployment Scaling
The Uni-Parser system is deployed on a cluster of 240 NVIDIA L40 GPUs (48 GB each), with 22 CPU cores and 90 GB of host memory per GPU. Under this setting, Uni-Parser parses over 16 million documents within 6 days in the fast mode. Moreover, the system scales to a hybrid cluster of up to 1,000 GPUs (A800 + L20), exhibiting near-linear throughput scaling with respect to available computational resources. These results indicate that Uni-Parser provides a computationally efficient, scalable, and stable infrastructure for large-scale document parsing.
@@ 4.1 Uni-Parser Data Engine
Step 2: Active Learning on Real Documents. After endowing the model with fundamental capabilities using large-scale synthetic data, we further improve its performance by iteratively incorporating real-world documents through an active learning framework. Our in-house training corpus is curated from multiple high-quality sources, including:
  • approximately 170 million pages of scientific literature spanning a broad range of disciplines;
  • over 140 million pages of patent documents collected from patent offices worldwide;
  • around 20 million pages of books, technical reports, and miscellaneous documents obtained from Fine-PDFs [12].
An active learning data flywheel, similar to the one used in MolParser [9], is employed to select the most informative samples for annotation. Specifically, cross-model inference consistency is used as the primary scoring criterion to estimate sample uncertainty and guide data selection during each active learning cycle. In addition, this process is integrated with a curriculum learning strategy, whereby the model is progressively trained on examples of increasing structural and semantic complexity, enabling stable optimization and improved generalization in challenging real-world scenarios.
@@ 5 Performance
Figure 9: Performance comparison of Uni-Parser with representative document parsing systems across two benchmarks. (a) Uni-Parser Benchmark, a benchmark proposed in this work to evaluate multimodal parsing capabilities on scientific documents and patent literature. (b) OmniDocBench-1.5 [29], a general-purpose benchmark for document parsing. Uni-Parser demonstrates strong and well-balanced performance across most evaluation dimensions, indicating favorable overall capability and robustness compared to existing methods.
5.1 Construction of the Uni-Parser Benchmark
The Uni-Parser benchmark is designed to evaluate layout detection and semantic content recognition in scientific documents. It comprises 150 PDF files collected from international patent documents and research articles, with an emphasis on diversity in document structure, subject matter, domain, and language. The distribution of the benchmark is summarized in Table 2. To ensure high data quality, the benchmark is annotated by more than ten domain experts, followed by two rounds of cross-review and an additional round of random quality inspection.
@@ 5.1 Construction of the Uni-Parser Benchmark
5.2 Overall Results
We evaluate scientific document parsing performance using an end-to-end evaluation protocol following OmniDocBench [29]. The Overall metric excludes molecular-related scores, as existing baselines generally lack molecular recognition and localization capabilities and therefore consistently obtain zero scores on these components. Table 3 reports the comparative results on the proposed Uni-Parser benchmark.
Table 3: Performance comparison of scientific document parsing methods on Uni-Parser Benchmark
Type Methods
Overall
excl. Mol.
$\uparrow$
Text
Edit
$\downarrow$
Form.
CDM
$\uparrow$
Tab.
TEDS
$\uparrow$
Tab.
TEDS-S
$\uparrow$
Read Order
Edit
$\downarrow$
Mol Loc
mAP@50
$\uparrow$
OCSR
Acc
$\uparrow$
Pipeline Tools MinerU2-pipeline 2.1.1 [43] 83.24 0.063 86.14 69.88 78.25 0.065 0.000 0.000
PP-StructureV3 [7] 82.41 0.069 91.42 62.71 75.21 0.040 0.000 0.000
Uni-Parser (Fast) 83.06 0.061 91.74 63.54 74.97 0.047 0.994 0.886
Uni-Parser (HQ) 89.74 0.022 92.37 79.04 82.79 0.018 0.994 0.886
Specialized VLMs MinerU2-VLM [43] 84.93 0.049 85.09 74.60 79.43 0.041 0.000 0.000
MinerU2.5 [43] 86.72 0.060 85.73 80.41 84.82 0.045 0.000 0.000
DeepSeek-OCR [46] 84.80 0.046 82.13 76.88 81.72 0.039 0.000 0.000
PaddleOCR-VL [6] 89.80 0.034 90.59 82.21 84.91 0.030 0.000 0.000
We further report results on OmniDocBench-1.5 [29], although this benchmark is not fully aligned with the core strengths of Uni-Parser. It mainly targets broad non-academic document domains (e.g., presentations and newspapers) and exhibits noticeable annotation noise. Nevertheless, for completeness, we include this evaluation. Leveraging the modular design of Uni-Parser, we replace the layout analysis module with PP-DocLayoutV2 [6] to accommodate this general-document OCR setting. In Table 4, this variant is denoted as Uni-Parser-G.
Table 4: Performance comparison of general document parsing methods on OmniDocBench-1.5 [29].
[…]
Pipeline Tools Marker-1.8.2 [42] 71.30 0.206 76.66 57.88 71.17 0.250
MinerU2-pipeline [43] 75.51 0.209 76.55 70.90 79.11 0.225
PP-StructureV3 [7] 86.73 0.073 85.79 81.68 89.48 0.073
Uni-Parser-G (Fast) 83.00 0.087 82.16 75.55 83.78 0.131
Uni-Parser-G (HQ) 89.75 0.047 88.83 85.12 89.07 0.099
General VLMs GPT-4o 75.02 0.217 79.70 67.07 76.09 0.148
InternVL3-76B [56] 80.33 0.131 83.42 70.64 77.74 0.113
InternVL3.5-241B [56] 82.67 0.142 87.23 75.00 81.28 0.125
Qwen2.5-VL-72B [1] 87.02 0.094 88.27 82.15 86.22 0.102
Gemini-2.5 Pro 88.03 0.075 85.82 85.71 90.29 0.097
Specialized VLMs Dolphin [10] 74.67 0.125 67.85 68.70 77.77 0.124
OCRFlux-3B [4] 74.82 0.193 68.03 75.75 80.23 0.202
Mistral OCR [23] 78.83 0.164 82.84 70.03 78.04 0.144
POINTS-Reader [17] 80.98 0.134 79.20 77.13 81.66 0.145
olmOCR-7B [33] 81.79 0.096 86.04 68.92 74.77 0.121
MinerU2-VLM [43] 85.56 0.078 80.95 83.54 87.66 0.086
Nanonets-OCR-s [19] 85.59 0.093 85.90 80.14 85.57 0.108
MonkeyOCR-pro-1.2B [15] 86.96 0.084 85.02 84.24 89.02 0.130
DeepSeek-OCR (9-crops) [46] 87.36 0.073 84.14 85.25 89.01 0.085
MonkeyOCR-3B [15] 87.13 0.075 87.45 81.39 85.92 0.129
dots.ocr [36] 88.41 0.048 83.22 86.78 90.62 0.053
MonkeyOCR-pro-3B [15] 88.85 0.075 87.25 86.78 90.63 0.128
MinerU2.5 [43] 90.67 0.047 88.46 88.22 92.38 0.044
PaddleOCR-VL [6] 92.86 0.035 91.22 90.89 94.76 0.043
Overall, the results indicate that Uni-Parser achieves a clear performance advantage in parsing scientific documents and patent literature. Moreover, owing to its modular architecture, Uni-Parser readily adapts to diverse document types while maintaining competitive accuracy. Notably, it operates with first-tier accuracy together with industrial-grade inference speed and stability, highlighting its suitability for large-scale real-world deployment.
@@ 5.3 Chemical Structure Parsing
We evaluate Uni-Parser on both the Uni-Parser Benchmark and the recent third-party BioVista benchmark [52]. As shown in Table 5(a), our group-based layout detection model, Uni-Parser-LD, outperforms our previous state-of-the-art molecular detection model, MolDet [9], while also establishing correspondences between molecules and their identifiers—a capability absent in MolDet. Table 5(b) shows that MolParser 1.5 consistently surpasses its predecessor across all benchmark subsets, delivering notable improvements in both accuracy and inference speed over mainstream open-source OCSR methods. Moreover, our approach remains highly competitive against costly commercial models [49, 48]. Representative case studies are provided in Appendix B.1.
@@
MolDet-Doc-L[9] 0.983 0.919
MolDet-General-L[9] 0.974 0.815
@@
MolDet-Doc-L[9] 0.961 0.871
MolDet-General-L[9] 0.945 0.815
BioMiner[52] 0.929 -
MolMiner[51] 0.899 -
@@
MolParser 1.0[9] 0.953 0.676 0.664 0.800
MolScribe[34] 0.617 0.274 0.168 0.417
@@
MolParser 1.0[9] 0.669 0.352 0.733 0.703
MolMiner[51] 0.774 0.497 0.185 0.507
MolScribe[34] 0.703 0.481 0.156 0.455
MolNexTR[5] 0.695 0.419 0.045 0.401
DECIMER[35] 0.545 0.326 0.000 0.298
@@
Table 6: Comparison of PDF parsing systems with OCSR support on a tiny test set, provided purely as a qualitative analysis rather than a quantitative evaluation.
@@
MathPix [20] 2025-08-27 100% 75.9% 59.6% - 66.1 s
MinerU.Chem [21] 2025-09-17 66.7% 63.1% 22.7% - $\sim$7 min